OpenStudy (anonymous):

Differentiate the following function with respect to x: (x+1)^3*(2x^2+3)*(x+3)^1/2 -------------------------- (x-4)^4 That's a fraction and the question hint says "Use logs".

OpenStudy (anonymous):

yeah i bet it does this is a nightmare

OpenStudy (anonymous):

\[f(x)=\frac{(x+1)^3(2x^2+3)(x+3)^{1/2}}{(x-4)^4}\]\[\ln(f(x))=3\ln(x+1)+\ln(2x^2+3)+\frac{1}{2}\ln(x+3)-4\ln(x-4)\]

OpenStudy (anonymous):

now taking the derivative is very easy of course you are taking the derivative of the log, so for your final answer you will have to multiply by the original function

OpenStudy (anonymous):

thank you so much